LOW-BARRIER HYDROGEN BONDING IN ELECTRONICALLY-EXCITED ACETYLACETONE

\underline{S. A. BROADBENT}, C. CHATTERJEE, S. M. WILSON, and P. H. VACCARO, Department of Chemistry, Yale University, P.O. Box 208107, New Haven, CT 06520; B. R. JOHNSON, Department of Chemistry, Rice University, P.O. Box 1892, Houston, TX 77251.

The cis-enol forms of β-diketones are stabilized by the action of strong intramolecular hydrogen bonding that mediates attendant proton-transfer processes. One of the simplest members of this class of compounds, acetylacetone (H_{3}C-CO-CH_{2}-CO-CH_{3}), has been demonstrated (both experimentally and theoretically) to possess an asymmetric equilibrium geometry with the two equivalent enol tautomers of the X̃^{1}A_{1} ground electronic state separated by a barrier of finite height. In contrast, recent ab initio calculations^{a} have suggested that the less-studied B̃^{1}B_{2} (π^{∗} ← π) manifold supports a symmetric (C_{2v}) minimum energy configuration. This excited-state structure would imply the presence of a low-barrier hydrogen bond (LBHB), in which the hydrogen is dislocated to a position equidistant from the two oxygen atom centers. Resonance Raman spectroscopy has been used to explore the prediction of an LBHB in acetylacetone and to elucidate the structural changes incurred by π^{∗} ←π electron promotion. Vapor-phase measurements were performed at five discrete excitation wavelengths, chosen to span the π^{∗} ← π absorption band (λ_{max} ≈ 262 nm) as well as pre-resonant and nonresonant portions of the spectrum. The acquired data displayed marked differences in their vibrationally-resolved intensity patterns. As the incident frequency was tuned toward resonance, vibrations involving distortion of the chelate ring were enhanced, whereas modes localized on the methyl groups did not exhibit significant enhancement. In resonant and preresonant spectra, numerous overtone and combination bands were observed in the 1620 − 2800 cm^{−1} and > 3000 cm^{−1} regions. These results are indicative of large geometry changes taking place upon π^{∗} ← π excitation, with theoretical analysis based on a time-dependent formalism for Raman scattering being utilized to quantify and interpret such structural modifications.

aH. P. Upadhyaya, A. Kumar, and P. D. Naik, J. Chem. Phys. \underline{110}, 11850 (2003).